Find the general solution of the following differential equation using the method of



undetermined coefficients: d^y/dx^2 - dy/dx + 2y = 2cosh2x

1
Expert's answer
2022-05-11T15:27:52-0400

At first, we solve the homogeneous equation: "g''-g'+2g=0". The characteristic equation has the form: "\lambda^2-\lambda+2=0". The solution is: "\lambda_1=\frac{1-i\sqrt{7}}{2}", "\lambda_2=\frac{1+i\sqrt{7}}{2}". Thus, "g=C_1e^{\lambda_1x}+C_2e^{\lambda_2x}"with "C_1,C_2\in{\mathbb{C}}" is the solution of equation: "g''-g'+2g=0". The solution can be rewritten in the following form: "g=k_1e^{\frac12x}\sin(\frac{\sqrt{7}}2x)+k_2e^{\frac12x}\cos(\frac{\sqrt{7}}2x)" with "k_1,k_2\in{\mathbb{C}}". We will try to find a solution of inhomogeneous equation in the form: "f=ae^{\alpha x}+be^{\beta x}" with "a,b,\alpha,\beta\in{\mathbb{C}}" . We receive: "f'=a\alpha e^{\alpha x}+b\beta e^{\beta x}", "f''=a\alpha^2e^{\alpha x}+b\beta^2e^{\beta x}". "f''-f'+2f=a(\alpha^2-\alpha+2)e^{\alpha x}+b(\beta^2-\beta+2)e^{\beta x}" . Right side of equation has the form: "e^{2x}+e^{-2x}". We put "\alpha=2", "\beta=-2" and receive: "f''-f'+2f=4ae^{2x}+8be^{-2x}=e^{2x}+e^{-2x}". From the latter we get: "a=\frac14", "b=\frac18". A solution is "f=\frac14e^{2x}+\frac18e^{-2x}". The general solution of the initial equation is: "y=f+g=k_1e^{\frac12x}\sin(\frac{\sqrt{7}}2x)+k_2e^{\frac12x}\cos(\frac{\sqrt{7}}2x)+\frac14e^{2x}+\frac18e^{-2x}" with "k_1,k_2\in{\mathbb{C}}".

Answer: the solution is: "y=k_1e^{\frac12x}\sin(\frac{\sqrt{7}}2x)+k_2e^{\frac12x}\cos(\frac{\sqrt{7}}2x)+\frac14e^{2x}+\frac18e^{-2x}" with "k_1,k_2\in{\mathbb{C}}".